OpenStudy (anonymous):

"Bobby has been planting blackberries in a garden in his backyard. Bobby started with blackberries in one square foot of garden modeled by the function s(x) = 30. He is allowed to convert more of the garden, and each additional square foot he maintains allows his blackberry plants to produce more by a rate of a(x) = 1.3^(x-1). Explain to Bobby how to create an equation to predict the number of blackberries he can expect based on the number of square feet he maintains. Describe how to determine the number of blackberries he will grow with 20 square feet. Use complete sentences." Help please! Ive been stuck on this question for 2 hours =c

OpenStudy (anonymous):

"Bobby has been planting blackberries in a garden in his backyard. Bobby started with blackberries in one square foot of garden modeled by the function s(x) = 30. He is allowed to convert more of the garden, and each additional square foot he maintains allows his blackberry plants to produce more by a rate of a(x) = 1.3^(x-1). Explain to Bobby how to create an equation to predict the number of blackberries he can expect based on the number of square feet he maintains. Describe how to determine the number of strawberries he will grow with 20 square feet. Use complete sentences."

OpenStudy (epoweritheta):

i dont understand this : Bobby started with blackberries in one square foot of garden modeled by the function s(x) = 30 ?? what does it mean ??

OpenStudy (anonymous):

im 100% lost with this question, i don't even understand how to start it

OpenStudy (anonymous):

@mathmale can you help? I'm not asking for an answer, im asking for help on how to actually do this ..

OpenStudy (anonymous):

@kennyj123 @SolomonZelman any ideas on how to start this? =c

OpenStudy (anonymous):

@hero any thougts?..

OpenStudy (anonymous):

I Googled the question and all I found was the following quote: 's(x) = 30 is a constant not a function.' Sorry I can't be of more help.

OpenStudy (anonymous):

I've been searching through my notes and I can't find anything..

OpenStudy (anonymous):

I'm going to move on and come back to this question later =c

hero (hero):

Actually, I just figured it out.

OpenStudy (anonymous):

You did? Please explain!

hero (hero):

s(x) = 30 is the the number of blackberries produced in one square feet a(x) = 1.3^(x - 1) is the rate of blackberries produced per x square feet f(x) = s(x)a(x) is the number of blackberries produced per x square feet f(x) = (30)(1.3^(x - 1))

hero (hero):

You have to find the number of blackberries produced with 20 square feet of garden space. f(20) = 30(1.3^(20 - 1))

OpenStudy (anonymous):

That...actually makes perfect sense to me o.o...

OpenStudy (anonymous):

HERO TRULY IS A HERO T^T THANK YOU SO MUCH . I've been stressing over this for so long i thought I'd never find out how to do it Dx

OpenStudy (anonymous):

and that actually helps me out with another one too. Thank you very very very much good sir/madam and have a wonderful day. I cannot express how thankful I am to you

hero (hero):

You're welcome